OpenStudy (anonymous):

WILL FAN AND MEDAL I need help with the equation 22=-5(3k+1) I have tried solving for k but I cant figure it out

Mehek (mehek14):

do you know how to distribute?

OpenStudy (anonymous):

YES

Mehek (mehek14):

then distribute -5(3k+1)

OpenStudy (anonymous):

IT BE 15K-5

OpenStudy (♪chibiterasu):

Now isolate K.

Mehek (mehek14):

-5 * 3k = -15

Mehek (mehek14):

-15k*

Mehek (mehek14):

you forgot the negative sign

OpenStudy (anonymous):

WHAT!!

OpenStudy (anonymous):

22=-15k-5 27=-15k

OpenStudy (anonymous):

like dat foh

Mehek (mehek14):

\(\bf{22 = -15k - 5}\) add 22 + 5 = 27 \(\bf{27=-15k}\) now divide

OpenStudy (anonymous):

HOW DO YOU ISOLATE THE VARIABLE

OpenStudy (anonymous):

divide

Mehek (mehek14):

move everything to one side and keep the variable on the other

OpenStudy (♪chibiterasu):

You have to cancel out the values on one side. To remove the -5 you add 5 to both sides which will not change the value of the equation.

OpenStudy (♪chibiterasu):

Doing this gives you the value of the unknown variable.

Mehek (mehek14):

since -15k = -15 x k you will divide 27/-15 and -15 x k / -15 that cancels out -15 on the right side leaving you with 27/-15 = k divide that to get your answer

OpenStudy (anonymous):

SO ITS -1.8

Mehek (mehek14):

yes

Mehek (mehek14):

k = -1.8

OpenStudy (♪chibiterasu):

Yeah. And if you want to make sure then you can input -1.8 into the equation and solve.

OpenStudy (anonymous):

i WISH I COULD GIVE YOU ALL MEDALS

OpenStudy (anonymous):

OOOOHHHH I FORGOT THANK YOU

Mehek (mehek14):

yw ^_^

OpenStudy (anonymous):

Mehek14 go to another question in algebra so i can give you a medal

OpenStudy (♪chibiterasu):

ARe you sure it isn't 32 and not 22?

OpenStudy (♪chibiterasu):

Because when I plug in the value I get 32.

OpenStudy (anonymous):

its 22 im sure i have the paper right here

OpenStudy (♪chibiterasu):

22 = -5(5.4 + 1) 22 = -5(6.4) 22 = -32

Mehek (mehek14):

you did it wrong chibiterasu :P

OpenStudy (♪chibiterasu):

I'm so stupid.

OpenStudy (♪chibiterasu):

Yeah thats right.

Mehek (mehek14):

\(\bf{22=-5(3*-1.8+1)\\22=-5(-5.4+1)\\22=-5(-4.4)\\22=22}\)

OpenStudy (♪chibiterasu):

Yeah I got 6.4 and not 4.4 because I mixed up the value sign. WOOPS